# Mekelweg 2, 2628 CD Delft, The Netherlands

### v1

@@ Abstract
The growing demand for natural interactions with technology underscores the importance of achieving realistic touch sensations in digital environments. Realizing this goal highly depends on comprehensive databases of finger-surface interactions, which need further development. Here, we present SENS3, an extensive open-access repository of multisensory data acquired from fifty surfaces when two participants explored them with their fingertips through static contact, pressing, tapping, and sliding. SENS3 encompasses high-fidelity visual, audio, and haptic information recorded during these interactions, including videos, sounds, contact forces, torques, positions, accelerations, skin temperature, heat flux, and surface photographs. Additionally, it incorporates thirteen participants’ psychophysical sensation ratings while exploring these surfaces freely. We anticipate that SENS3 will be valuable for advancing multisensory texture rendering, user experience development, and touch sensing in robotics.
@@
Texture Dataset Haptic Multisensory Sensation
@@ 1 Introduction
Recent trends in interactive system development emphasize a naturalistic approach to replicating human engagement with their physical surroundings in digital environments, promising enhanced user experience, accessibility, and improved digital communication [9]. While everyday tasks in the physical world involve engaging with objects through multiple senses [14], transferring this rich sensory information to the digital domain remains a challenge despite advancements in hardware and algorithms for capturing and recreating real-world sensory experiences [25].
Besides device and algorithm design, achieving naturalistic human-technology interactions relies on one more essential element: data. Thanks to the advancements in camera and microphone technologies, one can effortlessly capture and share their surrounding’s audio and visual information. Unfortunately, the same practicality does not apply to tactile data. When humans touch an object, they feel a rich array of tactile cues revealing its distinct surface properties, such as friction, roughness, thermal, and compliance [17]. High-fidelity data collection for each property requires specialized expertise and recording technology. This situation makes it challenging for regular users to record the tactile feel of every encountered surface instantly. A fundamental issue is that tactile cues, unlike light and sound waves, require physical contact. The resulting skin deformations depend on finger and surface properties and applied normal force and speed [18]; these can vary substantially even for the same user and surface, hugely influencing the recorded tactile data. Due to these reasons, only a few available databases [4, 22, 12, 7] include tactile recordings from interactions with surfaces. Moreover, existing databases often concentrate on a single or limited number of sensory cues insufficient for lifelike digitization of surfaces or neglect bare-finger interactions, the most natural way of interacting with our surroundings [15]. Finally, these databases mainly do not include psychophysical sensations humans perceive upon interaction; providing such information can tremendously help understand human texture perception or generate algorithms for machine perception.
@@
Dataset Visual Audio Tactile Exploration Interaction Sensations
[…]
@@
Sliding Tool tip -
CBSMR [21]
Image
@@
3D accelerations Sliding
Bare-finger
& tool-tip
@@
2D contact forces Lateral motion
@@
Sliding Bare-finger -
Learn2feel [18, 24] Image -
@@
Normal force
& geometric
measurements
Pressing Tool-tip -
Concurrent* [7]
Image
& video
Yes
@@
Sliding Bare-finger -
Skin temperature, heat flux, 3D contact forces & torques Static contact
SENS3 Image & video Yes 3D contact forces & torques, indentation depth Pressing Bare-finger Adjective ratings
3D contact forces & accelerations Tapping
3D contact forces, torques & accelerations, finger position & speed Sliding
@@ 2 Existing Haptic Texture Databases
Several databases have been made to date for digitizing tactile textures (see Table 1). The first comprehensive one, the Penn Haptic Texture Toolkit (HaTT), featured unconstrained tool-surface interactions with 100 surfaces, their models for haptic rendering, and high-resolution images of each surface [4]. The tactile data included contact forces, accelerations, and scan speed and was recorded with a custom-designed tool. The authors later demonstrated that the recorded contact accelerations could be used to re-create the vibratory feels of surfaces via a voice-coil actuator [6].
HaTT mainly focused on roughness and friction modalities, overlooking capturing multisensory aspects of texture perception [14]. To address this gap, Strese et al. developed two texture databases [21, 22]. They recorded data using multiple custom recording devices while exploring more than 100 surfaces via sliding, tapping, contour following, static contact, pressing, and holding. Their recording devices allowed mainly tool- or sensor-surface interactions. The databases contained contact accelerations, forces, thermal conductivity, reflectance, mass, volume, and interaction sounds and videos; check Table 1 for detailed distribution of collected data in each database. Similarly, Gao et al. created "OBJECTFOLDER2.0" [8], which incorporated tactile data (material, shape, texture, surface) along with audio and video of the tool interactions. Later, the authors demonstrated that these databases could be successfully used for texture classification via machine-learning techniques.
@@
Finally, only some of the databases [24, 18, 10] include user perception along with the physical data, often focusing on pairwise similarities. While these similarities provide insights into how interaction data shapes human perception, articulating precise statements on how changes in contact data correlate with sensations remains challenging and necessitates adjective ratings for each sensation [18].
@@ 3 SENS3 Database
We aimed to overcome the limitations of previous databases summarized in Section 2 with SENS3. We recorded the data while participants explored the surfaces with their bare fingers. As human fingertips show variety in geometry and mechanical properties [16, 19], we collected data from multiple people. Moreover, as previous studies [14] showed evidence that all tactile, visual, and auditory cues contribute to surface perception, we recorded multisensory information that spans these three senses. Additionally, we collected user sensation ratings to complement the physical recordings to understand human texture and machine perception and develop algorithms for effective texture rendering.
@@
Because human fingertips are soft and show nonlinear behavior with applied pressure and moving speed, the interaction data also vary as a function of these parameters [23]. Therefore, we collected contact forces and torques for all interactions. Due to the same reason, the data for the sliding experiments covered a spectrum of speeds (ranging from 0 to 200 mm/s) and applied forces (ranging from 0 to 1 N). Nonetheless, the rest of the interactions covered only one or two selected pressure ranges. In addition, we collected finger vibrations for active dynamic interactions (tapping and sliding), indentation depth for pressing, heat flux, and skin/surface temperature for static contact. See Table 1 for a detailed list of collected data in SENS3.
@@ 3.1 Selected surfaces & data collection apparatus
We built a custom data collection apparatus (see Supplementary Video), which was placed on an aluminum optical breadboard table (MB60120/M, Thorlabs) and mounted onto a robust supporting frame (PFM52502, Thorlabs). During the experiments, a monitor (UltraSharp 24, Dell) displayed graphical user interfaces (GUI).
@@ 3.2 Data collection procedure
Finger-surface interaction data: Two males, the first two authors of this paper, with an age of 26 years, participated in the finger-surface interaction data collection. Both participants gave informed consent and agreed to share their finger-interaction data publicly. Before each measurement, they washed their hands with soap and dried them using a towel. Then, each participant underwent an instruction and training process. They sat on a chair and put their arm on the armrest, facing the monitor. Afterward, their finger-surface interactions were collected while they explored the surfaces by applying static contact, pressing, tapping, and sliding. The detailed procedures for each measurement are described in the following, and Supplementary Video 1 visualizes each interaction.
@@
Following a similar approach to [3], we first measured the initial temperature of a surface sample via an infrared thermometer. Afterward, the thermistor was attached to the finger pad of the participant, whereas the 3D-printed frame with the heat flux sensor was positioned on the sample surface and then stabilized with weights. The measurement started with a sound cue while the participants’ fingers lay just over the samples, enabling the thermistor to measure the initial finger temperature. After hearing a sound cue, the participant was instructed to place their index and middle finger on the frame. To keep the contact area constant, they maintained a constant contact force of 3 N for 60 seconds for each surface.
@@
Perceptual data: Through the perceptual measurements, we aimed to record haptic psychophysical sensations felt by participants during the free multisensory exploration of surfaces. For this, we used the semantic differential method [17] with 15 points, where participants rated the surfaces based on eight opposing adjectives by moving sliders (implemented in MATLAB’s GUI) for each adjective. We selected a subset of surface adjectives used by [1]: rough-smooth, flat-bumpy, sticky-slippery, hot-cold, regular-irregular, fine-coarse, hard-soft, and wet-dry. Before the experiment, we asked the participants to wash and dry their fingers first and then explore all the surfaces without any sensory restrictions—they were free to interact with surfaces by seeing, hearing, and touching—to adjust their adjective limits mentally. Following this, we conducted mock trials with five random surfaces to acclimate the participants to the study. We randomly presented the surfaces to the participants during the experiment. We asked participants to explore the surface freely using all their three senses (visual, audio, and haptic) for fifteen seconds after they heard the sound cue. Although they could interact with the surfaces as they wished, we encouraged them to use all haptic exploratory procedures. After the exploration, the participants rated the surface feels by adjusting the scales in the GUI. The experimental procedure was conducted in accordance with the Declaration of Helsinki and approved by TU Delft’s ethics committee with the application number 3469. Three women and ten men with an average age of 26.84 years (standard deviation, SD: 2.034) participated in the experiment. All participants gave informed consent.
@@ 3.3 Collected data
SENS3.net hosts our database [^2] The empty download links will be directed to the 4TU research data repository upon acceptance.. The website was designed to provide a comprehensive resource for the recorded data and the recording procedures. It consists of three pages:
@@
    ‘Recording Setup’ page: includes an overview of the recording setup, data recording procedures, hardware list, and a figure depicting the camera and microphone arrangement.
@@
    ‘Surfaces’ page: hosts the collection of finger-surface interaction data organized into ten distinct categories. Users can click on a category to view the photos of each surface within it and download the associated data. The surface data are further aggregated based on different participants. The ‘Surfaces’ page also includes a ‘Metadata’ section, offering details about available files along with their descriptions and further information about surface physical properties.
@@ 4 Discussion
We presented SENS3, an open-source multisensory database of finger-surface interactions and corresponding psychophysical sensations. This database contains recordings of when two participants interacted with 50 surfaces by static contact, pressing, tapping, and sliding. Our database captures a wide range of sensory information, such as surface photographs, video and sounds of the interactions, contact forces and torques, finger accelerations and positions, skin temperature, and heat flux transfer. In addition to the physical recordings, we captured the psychophysical sensations generated during free exploration through adjective ratings given by thirteen participants. We hope that SENS3 will provide the necessary data for content design for multisensory user interfaces, understanding human touch, and providing robots with humanlike touch sensations.
In the future, we aim to make SENS3.net an open-source platform intended to be a collaborative hub where laboratories, companies, or designers from diverse backgrounds add their new collected data or texture rendering models designed for various interfaces, such as touch-based devices and wearables. This collective effort will accelerate progress in the field of multisensory user interfaces but also empower a wide range of applications, from enhancing accessibility for individuals with sensory impairments to revolutionizing entertainment experiences. We also plan to add measurements of surface characteristics and finger properties.

### v2

@@ Abstract
The growing demand for natural interactions with technology underscores the importance of achieving realistic touch sensations in digital environments. Realizing this goal highly depends on comprehensive databases of finger-surface interactions, which need further development. Here, we present SENS3www.sens3.net— an extensive open-access repository of multisensory data acquired from fifty surfaces when two participants explored them with their fingertips through static contact, pressing, tapping, and sliding. SENS3 encompasses high-fidelity visual, audio, and haptic information recorded during these interactions, including videos, sounds, contact forces, torques, positions, accelerations, skin temperature, heat flux, and surface photographs. Additionally, it incorporates thirteen participants’ psychophysical sensation ratings (rough–smooth, flat–bumpy, sticky–slippery, hot–cold, regular–irregular, fine–coarse, hard–soft, and wet–dry) while exploring these surfaces freely. Designed with an open-ended framework, SENS3 has the potential to be expanded with additional textures and participants. We anticipate that SENS3 will be valuable for advancing multisensory texture rendering, user experience development, and touch sensing in robotics.
@@
Surface Dataset Haptic Multisensory Sensation
@@ 1 Introduction
Recent trends in interactive system development emphasize a naturalistic approach to replicating human engagement with their physical surroundings in digital environments, promising enhanced user experience, accessibility, and improved digital communication [11]. While everyday tasks in the physical world involve engaging with objects through multiple senses [16], transferring this rich sensory information to the digital domain remains a challenge despite advancements in hardware and algorithms for capturing and recreating real-world sensory experiences [28] and understanding multisensory integration [5].
Besides device and algorithm design, achieving naturalistic human-technology interactions relies on one more essential element: data. Thanks to the advancements in camera and microphone technologies, one can effortlessly capture and share their surrounding’s audio and visual information. Unfortunately, the same practicality does not apply to tactile data. When humans touch an object, they feel a rich array of tactile cues revealing its distinct surface properties, such as friction, roughness, thermal, and compliance [20], depending on the applied exploratory procedures, e.g., sliding, static contact, and pressing [15]. High-fidelity data collection for each property and exploratory procedure requires specialized expertise and recording technology. This situation makes it challenging for regular users to record the tactile feel of every encountered surface instantly. A fundamental issue is that tactile cues, unlike light and sound waves, require physical contact. The resulting skin deformations depend on finger and surface properties and applied normal force and speed [21]; these can vary substantially even for the same user and surface, hugely influencing the recorded tactile data [26]. Due to these reasons, only a few available databases [6, 24, 25, 21] include tactile recordings from interactions with surfaces. Moreover, existing databases often concentrate on single or dual sensory cues, e.g., vision and tactile, audition and tactile, or tactile only, insufficient for lifelike digitization of surfaces or neglect bare-finger interactions, the most natural way of interacting with our surroundings [17]. Finally, these databases mainly do not include psychophysical sensations humans perceive upon interaction such as rough–smooth, flat–bumpy, sticky–slippery, hot–cold, regular–irregular, fine–coarse, hard–soft, and wet–dry; providing such information can tremendously help understand human texture perception or generate algorithms for machine perception.
@@
Dataset Audio/ Visual Tactile Exploration Interaction Sensations
Texture
count
Participant
count
HaTT
 [6]
Image
@@
Sliding Tool-tip No 100 1
CBSMR
 [24]
Audio, image
@@
108 1
3D accelerations Sliding Bare-finger & tool-tip
@@
No 184 1
2D contact forces
Lateral
motion
@@
Sliding Bare-finger No 120 1
Learn2feel
 [21]
Image
@@
Sliding Bare-finger No 10 1
Haptic
library [12]
Image No
Free
exploration
Bare-finger
Cluster
sorting
84 10
Skin temperature,
heat flux, 3D contact
forces & torques
Static
contact
SENS3
Audio, image
& video
3D contact forces
& torques,
indentation depth
Pressing Bare-finger
Adjective
ratings
50 2/ 13
3D contact forces
& accelerations
Tapping
3D contact forces,
torques & accelerations
finger position & speed
Sliding
@@ 2 Existing Haptic Texture Databases
Several databases have been made to date for digitizing tactile textures; see Table 1 for a list of these prior work. The first comprehensive one, the Penn Haptic Texture Toolkit (HaTT), featured unconstrained tool-surface interactions with 100 surfaces, their models for haptic rendering, and high-resolution images of each surface [6]. The tactile data included contact forces, accelerations, and scan speed and was recorded with a custom-designed tool. The authors later demonstrated that the recorded contact accelerations could be used to re-create the vibratory feels of surfaces via a voice-coil actuator [8].
HaTT mainly focused on roughness and friction modalities, overlooking capturing multisensory aspects of texture perception [16]. To address this gap, Strese et al. developed two texture databases [24, 25]. They recorded data using multiple custom recording devices while exploring more than 100 surfaces via sliding, tapping, contour following, static contact, pressing, and holding. Their recording devices allowed mainly tool- or sensor-surface interactions. The databases contained contact accelerations, forces, thermal conductivity, reflectance, mass, volume, and interaction sounds and videos; check Table 1 for detailed distribution of collected data in each database.
@@
Finally, only some of the databases [27, 21, 12] include perception data along with the physical one, often focusing on pairwise similarities. While these similarities provide insights into how interaction data shapes human perception, articulating precise statements on how changes in contact data correlate with sensations remains challenging and necessitates adjective ratings for each sensation [21].
@@ 3 SENS3 Database
We aimed to overcome the limitations of previous databases summarized in Section 2 with SENS3. We recorded the physical interaction data while participants explored the surfaces with their bare fingers. As human fingertips show variety in geometry and mechanical properties [18, 22], we collected this data from two people. Moreover, as previous studies [16] showed evidence that all tactile, visual, and auditory cues contribute to surface perception, we recorded multisensory information that spans these three senses. Additionally, we collected sensation ratings from 13 people to complement the physical recordings to understand human texture and machine perception and develop algorithms for effective texture rendering.
@@
Because human fingertips are soft and show nonlinear behavior with applied pressure and moving speed, the interaction data also vary as a function of these parameters [26]. Therefore, we collected contact forces and torques for all interactions. Due to the same reason, the data for the sliding experiments covered a spectrum of speeds (ranging from 0 to 200 mm/s) and applied forces (ranging from 0 to 1 N). Nonetheless, the rest of the interactions covered only one or two selected pressure ranges. In addition, we collected finger vibrations for active dynamic interactions (tapping and sliding), indentation depth for pressing, heat flux, and skin/surface temperature for static contact. See Table 1 for a detailed list of collected data in prior databases and SENS3.
@@ 3.1 Selected surfaces & data collection apparatus
3.2 Data collection apparatus
Figure 2: Data recording apparatus: 1. Cameras, 2. data acquisition board, 3. heat flux sensor, 4. microphone, 5. linear stage, 6. position sensor, 7. armrest, 8. force sensor, 9. accelerometer, 10. selection of surfaces.
We built a custom data collection apparatus (see Fig. 2 and Supplementary Video), which was placed on an aluminum optical breadboard table (MB60120/M, Thorlabs) and mounted onto a robust supporting frame (PFM52502, Thorlabs). During the experiments, a monitor (UltraSharp 24, Dell) displayed graphical user interfaces (GUI).
@@ 3.3 Data collection procedure
Finger-surface interaction data: Two males, the first two authors of this paper, with an age of 26 years, participated in the finger-surface interaction data collection.
Both participants gave informed consent and agreed to share their finger-interaction data publicly. Before each measurement, they washed their hands with soap and dried them using a towel. Then, each participant underwent an instruction and training process. They sat on a chair and put their arm on the armrest, facing the monitor. Afterward, their finger-surface interactions were collected while they explored the surfaces by applying static contact, pressing, tapping, and sliding. The detailed procedures for each measurement are described in the following, and Supplementary Video 1 visualizes each interaction. Recording all types of physical interaction data from one participant for fifty surfaces took approximately (not consecutively) 24 hours.
@@
Following a similar approach to [4], we first measured the initial temperature of a surface sample via an infrared thermometer. Afterward, the thermistor was attached to the participant’s index finger pad, whereas the 3D-printed frame with the heat flux sensor was positioned on the sample surface and then stabilized with weights. The measurement started with a sound cue while the participants’ fingers lay just over the samples, enabling the thermistor to measure the initial finger temperature. After hearing a sound cue, the participant was instructed to place their index and middle finger on the frame. They maintained a constant contact force of 3 N for 60 seconds for each surface with both fingers to keep the contact area constant.
@@
Perceptual data: Through the perceptual measurements, we aimed to record haptic psychophysical sensations felt by participants during the free multisensory exploration of surfaces using their dominant index finger. For this, we used the semantic differential method [20] with 15 points, where participants rated the surfaces based on eight opposing adjectives by moving sliders (implemented in MATLAB’s GUI) for each adjective. We selected a subset of surface adjectives used by [1]: rough–smooth, flat–bumpy, sticky–slippery, hot–cold, regular–irregular, fine–coarse, hard–soft, and wet–dry. Before the experiment, we asked the participants to wash and dry their fingers first and then explore all the surfaces without any sensory restrictions—they were free to interact with surfaces by seeing, hearing, and touching—to adjust their adjective limits mentally. Following this, we conducted mock trials with five random surfaces to acclimate the participants to the study. During the experiments, we randomly presented the surfaces to the participants. We asked participants to explore the surface freely using all their three senses (visual, audio, and haptic) for fifteen seconds after they heard the sound cue. Although they could interact with the surfaces as they wished, we encouraged them to use all haptic exploratory procedures explained during mock trials. After the exploration, the participants rated the surface feels by adjusting the scales in the GUI. The experimental procedure followed the Declaration of Helsinki and was approved by TU Delft’s ethics committee with application number 3469. Three women and ten men with an average age of 26.84 years (standard deviation, SD: 2.034) participated in the experiment. All participants gave informed consent. The perceptual data collection procedure took 2 hours per participant.
@@ 3.4 Collected data
SENS3.net hosts our database. The website was designed to provide a comprehensive resource for the recorded data and the recording procedures. It consists of three pages:
@@
    ‘Recording Setup’ page: includes an overview of the recording setup, data recording procedures, GUI, hardware list, and a figure depicting the camera and microphone arrangement.
@@
    ‘Surfaces’ page: hosts the collection of finger-surface interaction data organized into ten distinct categories. Users can click on a category to view the photos of each surface within it and download the associated data. The surface data are further aggregated based on different participants. The ‘Surfaces’ page also includes a ‘Metadata’ section, offering details about available files along with their descriptions and further information about surface physical properties; see Table 2).
  4. 4.
    ‘About’ page: contains citation detail for the paper and additional information about the authors.
Table 2: Metadata for the recorded data files. (num) represents the allocated number tag of the surface.
File Description
Tapping
forces.csv
contact forces and torques in 6 columns
- Fx, Fy, Fz, Tx, Ty, Tz [N and Nmm]
accelerations.csv
accelerations in 3 columns
-Ax, Ay and Az [g]
audio.wav audio data
video(num)_1.mp4 side view video
video(num)_2.mp4 top view video
Pressing
forces.csv
contact forces and torques in 6 columns
- Fx, Fy, Fz, Tx, Ty, Tz [N and Nmm]
stageposition.csv position of the linear stage [mm]
video(num)_1_2mms.mp4 side view video
video(num)_2_2mms.mp4 top view video
Static Contact
forces.csv
contact forces and torques in 6 columns
- Fx, Fy, Fz, Tx, Ty, Tz [N and Nmm]
temperature.csv index finger temperature [°C]
material_temperature.csv surface temperature of the materials [°C]
heatflux.csv
heat flux between middle
finger and material [W/m2]
Sliding
Material_sensor_(num).csv
contact forces and torques in 6 columns
- Fx, Fy, Fz, Tx, Ty, Tz [N and Nmm],
accelerations in 3 columns,
Ax, Ay and Az [g]
Material_IR_pos_(num).csv
4 columns of data - elapsed time [s],
speed [mm/s], position X [$m$],
position Y [$m$]
Material_(num).mp4 top view video
Material_(num).wav audio data
Sensation rating
Order_par(num).csv order of material presented to participant
Ratings_par(num).csv adjective ratings
Video(num).mp4 isometric view video
material(num).tif top view image
thickness.csv thickness of the materials
@@
4 Results of the perceptual experiments
[…]
Table 3: Rotated Component Matrix
Principal Components
Adjectives 1 2 3 4
Rough–Smooth -0.7878 -0.0689 0.1380 0.5593
Flat–Bumpy 0.9263 0.2261 0.1576 -0.1397
Sticky–Slippery 0.0183 -0.0629 0.9605 -0.1643
Warm–Cold -0.2606 -0.4775 -0.1735 0.7923
Regular–Irregular 0.9246 0.1174 -0.0213 -0.1153
Fine–Coarse 0.9237 0.1741 0.0140 -0.3093
Hard–Soft 0.2604 0.8683 -0.1229 -0.3898
Wet–Dry 0.2543 0.2105 0.1923 -0.8912
Table. 3 shows that the first rotated component, accounting for 66% of the variance, encompasses adjective ratings such as "rough–smooth," "flat–bumpy," "regular–irregular," and "fine–coarse," indicating roughness cue. Similarly, the second rotated component, explaining 16% of the total variance, is characterized by the "hard–soft" adjective rating, signaling compliance cue. The third rotated component, describing 9% of the total variance, is represented by the "sticky–slippery" adjective rating, indicating the friction cue. Lastly, the fourth rotated component, capturing 4% of the total variance, is associated with adjective ratings such as "warm–cold" and "wet–dry," hinting at thermal cues.
@@ 5 Discussion
We presented SENS3, an open-source multisensory database of finger-surface interactions and corresponding psychophysical sensations. This database contains recordings of when two participants interacted with 50 surfaces by static contact, pressing, tapping, and sliding. Our database captures a wide range of sensory information, such as surface photographs, video and sounds of the interactions, contact forces and torques, finger accelerations and positions, skin temperature, and heat flux transfer. In addition to the physical recordings, we captured the psychophysical sensations generated during free exploration through adjective ratings given by thirteen participants. We hope that SENS3 will provide the necessary data for content design for multisensory user interfaces, understanding human touch and multisensory integration, and providing robots with humanlike touch sensations.
Despite our diligent efforts, the current dataset version exhibits some limitations. It comprises finger-surface interaction recordings from two trained male participants and perceptual ratings from thirteen participants, all collected by exploring fifty surfaces. This limited participant pool hampers our ability to capture the full spectrum of finger biomechanics and human perception. Furthermore, the surface diversity within the dataset represents merely a subset of surfaces encountered in daily life. Additionally, separating physical interaction recordings and adjective ratings poses a challenge in establishing direct correlations between sensations and physical data. Given the interconnected nature of adjective ratings and the inability of a single exploratory method to encapsulate all these sensations, we opted for free exploration, integrating all four exploratory procedures, to gather comprehensive perceptual data [19, 3] and high-quality finger-surface interaction data separately with distinct exploratory procedure.
We designed SENS3 as open-ended so that we can address the shortcomings mentioned above. We envision augmenting the dataset with data collected from a more diverse range of surfaces and participants, encompassing variations in age, gender, and finger biomechanics. Our expansion plans also incorporate surface characteristics, finger biomechanical measurements, and simultaneous perceptual data collection separately during each exploratory procedure.
We aim to make SENS3.net an open-source platform intended to be a collaborative hub where laboratories, companies, or designers from diverse backgrounds add their new collected data or texture rendering models designed for various interfaces, such as touch-based devices and wearables. This collective effort will accelerate progress in the field of multisensory user interfaces but also empower a wide range of applications, from enhancing accessibility for individuals with sensory impairments to revolutionizing entertainment experiences.